OpenStudy (anonymous):

26. The periodic time of a mass suspended by a spring (force constant k) is T. If the spring is cut into three equal parts, what will be the force constant of each part. If the same mass be suspended from one part, what will be the periodic time.

OpenStudy (anonymous):

@Vincent-Lyon.Fr

OpenStudy (anonymous):

if the spring is cut into three equal parts.. what do you think happens to the force constant?

OpenStudy (anonymous):

it remains the same

OpenStudy (anonymous):

Answer this when two springs of spring constants K are connected in series.. what happens to the spring constant of the resulting spring?

OpenStudy (anonymous):

1/k = 1/k1 + 1/k2

OpenStudy (anonymous):

correct.. so if the spring constants are same then the resulting constant would be 1/Kr = 1/K + 1/K = 2/K hence Kr = K/2 so what you learnt?? if two springs of same spring constant are put together in series, the spring constant goes down by a factor that is equal to the number of springs put togther so now what happens if i CUT the spring into three equal halfs.. what do you think happens to the K of each part? ;-)

OpenStudy (anonymous):

It Becomes 3 times

OpenStudy (anonymous):

correct.. so now spring constant of each spring is 3 times what is the relation between Periodic time and spring constant?

OpenStudy (anonymous):

@Mashy

OpenStudy (anonymous):

yea yea m here don't worry :P

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

T = 2 pie sqrt(m/k)

OpenStudy (anonymous):

Ryt ?

OpenStudy (anonymous):

you should know the relation between T and k and m \[T =2 \pi \sqrt{ \frac{m}{k}}\]

OpenStudy (anonymous):

yup.. so .. take that as equation one now develop a new equation for NEW T .. (lets call it T prime) and the new spring constant (3k).. but the mass is the same (given in the question).. this be the equation 2 divide the two equations to get your T prime

OpenStudy (anonymous):

Thxxxx...u helped me a lot...

OpenStudy (anonymous):

no problem :D.. thats what we do ;-)!